Instructions for the preparation of a manuscript for the special issue ”Nanostructures in Photovoltaics”

Richard Corkish^{a} ^{∗}, Stefan Kettemann^{b}, and Jenny Nelson^{c}

aCenter for Theoretical Physics, University of New South Wales Sydney 2052 Australia

bI. Institut f. Theoretische Physik Jungiusstr. 9, 20355 Hamburg Germany

cPhysics Department Imperial College London, SW7 2BZ, United Kingdom

Abstract

These pages provide you with an example of the layout and style which we request you to adopt during the preparation of your paper. Please insert your abstract here.

PACS: please insert PACS-1995 classification codes here. Keywords: please insert 3-4 keywords here.

1. Introduction

These guidelines are specifically for manuscripts intended for the special issue ”Nanostructures in Photovoltaics”. If you are in doubt please contact nip01 or refer to the instructions for authors for Physica E [1].

2. Instructions for preparation of a manuscript in LaTEX

Original material: Submission of a manuscript implies that it is not being simultaneously considered for publication elsewhere and that the authors have obtained the necessary authority for publication.

Types of contributions: Original research papers [...] should contain an Abstract (of up to 200 words) and a Conclusions section, which, translates the results into terms readily accessible to most readers.

Manuscript preparation: All manuscripts should be written in proper English. For electronic submission, LaTEX files using the style file espcrc2.sty (available from this web site) are acceptable. Alternatively, word processor files may be submitted.

One Paper copy of a manuscript should be submitted in addition, to be forwarded by the editors to a referee.

Structure: Please adhere to the following order of presentation: Article title, Author(s), Affiliation(s), Abstract, PACS codes and keywords, Main text, Acknowledgements, Appendices, References, Figure captions, Tables.

Corresponding author: The name, complete postal address, telephone and fax numbers and the E-mail address of the corresponding author should be included with the submission.

PACS codes / keywords: Please supply one or more relevant PACS-1995 classification codes and 3-4 keywords of your own choice for indexing purposes.

∗Footnotes should appear on the first page only to indicate your present address (if different from your normal address), research grant, sponsoring agency, etc. .

References: References to other work should be consecutively numbered in the text using square brackets and listed by number in the Reference list.

Illustrations: With an electronic submission, Postscript figures are acceptable, to be included with the LaTEX manuscript in a uuencoded, gzipped tar file.

Colour plates: Figures may be published in colour, if this is judged essential by the editor. The publisher and the author will each bear part of the extra costs involved. Further information is available from the publisher.

Length: Please adhere to the length restriction detailed in the letter inviting you to contribute to this special issue. Physica E has about 675 words on a printed journal page. You may use the LaTEX style file espcrc2.sty to estimate the length of your manuscript.

3. Submission

Please submit your manuscript by September,

2001 to nip01@mpipks-dresden.mpg.de and the

hardcopy to

Mandy Stiegler Max Planck Institute for the Physics of Complex Systems Nöthnitzer Str. 38 01187 Dresden Germany.

4. After acceptance

Notification: You will be notified by the Editor of the journal of the acceptance of your article by October 30th.

Copyright transfer: In the course of the production process you will be asked to transfer the copyright of the article to the publisher. This transfer will ensure the widest possible dissemination of information.

Acknowledgements

This file is based on the latex file, as written and compsed by Bernhard Mehlig, for Dynamics of Complex Systems (”DYNCOMPS”) 1999. We are gratefull to Peter Fulde for his continous support and stimulating discussions, without which this workshop would not have been possible.

References

1. Instructions for authors for Physica E are available in electronic format under

                                                 http://www.elsevier.nl